OpenStudy (anonymous):

find the unit tangent and curvature for the space curve x=t-t^3/3, y=t^2 , z=t+t^3/3

OpenStudy (anonymous):

The unit tangent vector defines a function that calculates the tangent vector with a magnitude of 1 (a unit vector) of a particular vector function. The equation for the unit tangent vector is: \[T=\frac{V}{\left| V \right|}\] In this case: \[T=\frac{<t-\frac{t^3}{3},t^2,t+\frac{t^3}{3}>}{\sqrt{({t-\frac{t^3}{3})^2+(t^2)^2+(t+\frac{t^3}{3})^2}}}\] I'll type another post about the curvature

OpenStudy (anonymous):

ohk the whole problem carries 14marks

OpenStudy (anonymous):

thet^2/3 is a fraction not exponent

OpenStudy (anonymous):

The curvature is defined as the rate at which a curve changes direction, and can be defined multiple ways. \[k=\frac{dT}{ds}=\frac{1}{\left| V \right|}\frac{|dT|}{|dt|}=\frac{\left| A \times V \right|}{\left| V \right|}\] Where ; T=unit tangent vector s=position vector function V=velocity vector function A= acceleration vector function In most cases you should use the first form.

OpenStudy (anonymous):

ohkay... but in class when we looked for the curvature we took the tangent into tne quotient formular

OpenStudy (anonymous):

like now m confused y do we have to use the quotient rule when we got the curvaure formular

OpenStudy (turingtest):

hold on there, the unit tangent for a vector function \(\vec r(t)\) is\[\vec T={\vec r'(t)\over\|\vec r'(t)\|}\]so you gotta take the derivative first

OpenStudy (anonymous):

yep... i agree with turningtest so when we look for the tangent we first look for the derivative and the derivative of the magnitude

OpenStudy (turingtest):

the curvature is\[\kappa={\|\vec T'(t)\|\over\|\vec r'(t)\|}\]

OpenStudy (turingtest):

so once you have the unit tangent just take the derivative, find the magnitude of that at the point in question, and divide by the magnitude of the derivative of the parametric curve, which we already found

OpenStudy (anonymous):

The curvature is just the derivative the unit tangent vector divided by r'(t)=v(t). And yes, what turingtest said is the same thing as what I said r'(t)=v(t), I also don't know how to put the arrow on top.

OpenStudy (anonymous):

yeap... that is clear to me so what is the answer you get

OpenStudy (turingtest):

or okay, no point needed, I guess they just want the general formula

OpenStudy (turingtest):

@abstracted \vec{whatever} = \[\vec{whatever}\]

OpenStudy (turingtest):

but yours doesn't take the derivative is the problem

OpenStudy (anonymous):

guys ur making to get confused what is the answer u get on the unit tangent

OpenStudy (turingtest):

\[\vec T={\vec V'\over\|\vec V'\|}\]

OpenStudy (anonymous):

In the second equation of my first post, I did forget to actually differentiate it, so ignore that one. But, \[\vec{T}=\frac{\vec{r'(t)}}{|\vec{r'(t)}|}=\frac{\vec{v(t)}}{|\vec{v(t)}|}\]

OpenStudy (turingtest):

ok I just figured out you meant that, sorry for the confusion you just didn't say that at first, and I didn't read every post :P

OpenStudy (anonymous):

Yea, it would've been simpler if I said r'(t) instead, I just copied straight from my old book.

OpenStudy (anonymous):

i know the formulas bt can u just help me with the answers

OpenStudy (turingtest):

no, I'm afraid you have to participate, we can't just give out answers... during this time you should have found \(\|\vec r'(t)\|\) and \(\vec r'(t)\) if not, try that first; you need it in the formula we are writing

OpenStudy (anonymous):

the answer for the unit tangent i got [(1-t^2) + 2t +( 1+t^2)] / [root of 2 (t^2 +1)^2

OpenStudy (turingtest):

if I could find a pen I could check that :P

OpenStudy (turingtest):

careful with the square root in the magnitude

OpenStudy (turingtest):

\[\|\vec r(t)\|=\sqrt{(1-t^2)^2+(2t)^2+(1+t)^2}\]

OpenStudy (anonymous):

what about the magnitude

OpenStudy (turingtest):

\[\|\vec r(t)\|=\sqrt{1-2t^2+t^4+4t^2+1+2t^2+t^4}\]

OpenStudy (turingtest):

that is the magnitude

OpenStudy (anonymous):

yes i agree but we got same factors

OpenStudy (turingtest):

\[\|\vec r(t)\|=\sqrt{2+4t^2+2t^4}=\sqrt{2(1+t^2)^2}=\sqrt2(1+t^2)\]I have to go teach a class I'm afraid, I hope someone else can continue helping you good luck!

OpenStudy (anonymous):

ok